CH/Meet/95/03

Minutes

Charissa Group Meeting

Meeting held on: 24 Nov 1995

Meeting held at: Daresbury

Present:  All physicists including students
Distribution: charmeet@nnga.dl.ac.uk, by e-mail/file-server

Minutes

This Charissa Physics Meeting attended by all physics staff and students was preceded by a summary of recent developments by BRF and ended with a discussion of MEGHA progress. These items are minuted below. The Physics Meeting was also followed on 25 November by the first ever Analysis Meeting, attended by students and postdocs involved in the analysis of Charissa data.

Minutes: this meeting was informal and the minutes of the meeting of 11 October 1995 will be considered at the next meeting, along with these minutes.

1. Welcome: The new students Graham Dillon (York), Mark Shawcross (Surrey) and Colin Freeman (Birmingham) were welcomed.

2. Report on the ANU:

Progress at the ANU: the insurance claim for the beamline had been resolved and ANU were proceeding with repairs. The first quadrupole repair was complete. The beamline installation was underway and progressing well. The 14UD was operating well after a recent routine foil change and the end of the strike action.

Progress at the ANU: The Jan/Feb 1996 running is being coordinated by SPGC. BRF is coordinating a planned March/May period including MEGHA installation and commissioning (volunteers welcome). Running in July 1996 is planned to include the first physics use of MEGHA and will coincide with the dedication of the Linac hopefully (around second week of July).

3. Other Experiments:

Strasbourg: awaiting beam to test beamline. Plan a physics run as soon as possible thereafter. RPW has a new proposal for the next PAC.

GANIL: WNC, NC, MF and NMC returned this week from a run on SPEG. A boron 17 run is planned around Feb 1996. Two new proposals are being put to the PAC next week, with WNC as spokesperson for one of them. Preamps, amps and test equipment are being built.

4. Papers:

This year 12 papers have been published at last count, and another 4 are with the referees, and more in preparation. The number of PhD theses awarded this year is 5.

Some backlog exists still with ANU data, and should be cleared.

5. EPSRC Grant Spend:

A review of the spend was completed in the last week. Spare sheet and strip detectors will be purchased for MEGHA. Some funds will be held until March before final committment.

6. Preparations for new grant:

The request is due in March 1996, for 1996–2000. It will be a big job, with new rules. A report of around 20 pages for Charissa is required. It is noted that the recent EPSRC review gives good support for exploitation of the investment in ANU running. A profile of the likely themes of the case for support was suggested, and input was invited.

7. Presentations on analysis in progress:

These were given by SMS, NC, MF, SH, GRK, RC, SPGC, PF and RLM.

8. MEGHA update:

DLW outlined the status of the chamber and highlighted points requiring attention, namely:

i. rearrangement of the beamline after the chamber, to move the beam

dump nearer to the target as requested by the ANU, ii. beam collimation and antiscatter at the chamber entrance.

Possible solutions were discussed.

W.N. Catford, Secretary. November 25, 1995 (released November 25, 1995)